Find an equation of the line containing (4,3) and (12,8)
Answer by mananth(16949)

x1 y1 x2 y2
4 3 12 8

slope m = (y2-y1)/(x2-x1)
( 8 - 3 )/( 12 - 4
( 5 / 8 )
m= 5/ 8

Plug value of the slope and point ( 4, 3 ) in
Y = m x + b
3.00 = 2 1/2 + b
b= 3 - 2 1/2
b= 1/ 2
So the equation will be
Y = 5/8 x + 1/2
m.ananth@hotmail.ca